# Revisiting Acoustic Similarity in Emotional Speech and Music via Self-Supervised Representations

### v1

@@ I Introduction
Nevertheless, the shared acoustics cues between speech and music, particularly those encoded by SSL models, remain largely unexplored. While SSL-based models have demonstrated impressive performance in both tasks, the extent to which these models capture and utilize overlapping acoustic features across speech and music is not well understood. This gap limits progress in developing cross-domain insights and hinders the potential for improving SER and MER systems.
@@ III-A Dataset
We use the RAVDESS dataset [14], an audio-visual collection consisting of acted affective recordings of speech and song. The dataset features recordings of 12 female and 12 male professional actors speaking English with a North American accent. In our study, we focus on the audio recordings, excluding the visual input. These recordings contain only vocal performances without accompanying instrumental music, thereby excluding the instrumental influence on the acoustic similarity. Since the song recordings of one female actor are missing, we discard her speech recordings as well to ensure an equal amount of speech and song data for model training. The text content in both the speech and song recordings is semantically neutral and identical, allowing us to concentrate on the acoustic properties by eliminating potential effects caused by lexical content (e.g., prosody variations due to word pronunciation). These are the key reasons we chose RAVDESS, despite the availability of larger datasets.
@@ V-A2 Results and Discussions
For SER, it can be observed that performance on angry and fearful speech is generally the best, followed by calm, across all speech and music models. This suggests that both speech and music models favor the acoustic characteristics of these emotions in speech. However, for MER, recognizing fearful and calm proves to be more challenging, particularly in the case of fearful with MERT. In contrast, angry music becomes easier to recognize, even for speech models. Moreover, neutral performs much better on music with MERT than any other neutral cases. Additionally, angry music with W2V2 does not show a performance drop, unlike other emotions with W2V2 for either speech or music. We believe that the deep layers of W2V2 might be particularly effective at modeling acoustic characteristics related to angry music, such as intense rhythm. These observations suggest that different emotions in speech and music exhibit distinct acoustic characteristics, highlighting the emotion bias problem in both speech and music SSL models, and strengthening the findings of [10]. Finally, other emotions all have their respective variations on speech or music with speech models or the music model, but the overall patterns remains similar. This further verifies previous findings that shared acoustics exist between speech and music [6, 13].
@@ V-B1 Experimental Settings
For the baseline and WS models, we use the AdamW optimizer with a learning rate of $1e{-3}$. For the PEFT models, we use the AdamW optimizer with a learning rate of $1e{-4}$. For the same reason as layerwise probing, each model is trained for 300 epochs at each stage to achieve its best possible performance.
@@ V-B2 Results and Discussions
Table III summarizes the domain adaptation results using all three approaches and models. Note that the results of SER and MER from the same model are not comparable, as they are different tasks. We only compare SER with SER and MER with MER. It can be seen that for the speech models (W2V2 and HuBERT), fine-tuning on either domain in stage one always helps the other domain in stage two. For example, the performances of SER followed by MER (74.88, 80.30, 87.20 for W2V2; 84.24, 82.76, 93.10 for HuBERT) are always better than directly conducting SER (73.89, 79.31, 87.20 for W2V2; 83.25, 80.30, 88.18 for HuBERT). Furthermore, PEFT works extremely effectively for HuBERT, achieving the best performance (93.10) and the largest improvement (4.92). For the music model, while several performances are improved after domain adaptation (e.g., 72.77 vs. 61.85 for SER and 93.60 vs. 91.63), such a phenomenon does not generally hold. This finding further verifies our conclusion of finding III in Section V-A that speech SSL models are further trainable with music acoustics yet further training of music SSL models with speech acoustics may be less effective.
@@ V-C2 Results and Discussions
(I) In terms of distance ranking, all the speech and music models exhibit consistent patterns: angry and fearful have the smallest distances, followed by happy, sad, neutral, and calm. These consistent patterns indicate that the SSL models have a similar ability to capture acoustic similarities related to emotion in their respective representations. Since speech and music share the same text in RAVDESS, the low FAD scores for certain emotions indicate that their acoustics are largely similar between speech and music For example, angry speech and music are more acoustically similar than calm speech and music.
@@ VI Conclusion
As the first work exploring acoustic similarity between emotional speech and music through SSL models, we conduct layerwise probing to explore the models’ ability to capture emotion-relevant acoustic cues and uncover the commonality and differences between speech and music SSL models. Moreover, aiming to leverage each domain to enhance the emotion recognition performance for the other domain, we implement and compare three approaches, demonstrating efficacy of the PEFT approach for cross-domain adaptation and noticing the domain-specificity of the models. Finally, we employ FAD to measure the acoustic similarity between speech and music representations on each emotion, revealing different behaviors of SSL models and the emotion-bias issue in both domains. Our study provide new insights into the acoustic similarity between emotional speech and music, particularly with SSL models. We also demonstrate the potential of leverage knowledge from each domain to facilitate the other. As there are limited datasets suitable for further investigation and fair comparison, we acknowledge the limitations of our work and plan to collect new datasets for extending our study in the future.

### v2

# Exploring Acoustic Similarity in Emotional Speech and Music via Self-Supervised Representations

@@ I Introduction
Nevertheless, the shared acoustic cues between speech and music, particularly those encoded by SSL models, remain largely unexplored. While SSL-based models have demonstrated impressive performance in both tasks, the extent to which these models capture and utilize overlapping acoustic features across speech and music is not well understood. This gap limits progress in developing cross-domain insights and hinders the potential for improving SER and MER systems.
@@ III-A Dataset
We use the RAVDESS dataset [14], an audio-visual collection consisting of acted affective recordings of speech and song. The dataset features recordings of 12 female and 12 male professional actors speaking English with a North American accent. In our study, we focus on the audio recordings, excluding the visual input. These recordings contain only vocal performances without accompanying instrumental music, thereby excluding the instrumental influence on the acoustic similarity. Since the song recordings of one female actor are missing, we discard her speech recordings as well to ensure an equal amount of speech and song data for model training. The text content in both the speech and song recordings is semantically neutral and identical, allowing us to concentrate on the acoustic properties by eliminating potential effects caused by lexical content (e.g., prosody variations due to word pronunciation). These are the key reasons we have chosen RAVDESS, despite the availability of larger datasets.
@@ V-A2 Results and Discussions
For SER, it can be observed that performance on angry and fearful speech is generally the best, followed by calm, across all speech and music models. This suggests that both speech and music models favor the acoustic characteristics of these emotions in speech. However, for MER, recognizing fearful and calm proves to be more challenging, particularly in the case of fearful with MERT. In contrast, angry music becomes easier to recognize, even for speech models. Moreover, neutral performs much better on music with MERT than any other neutral cases. Additionally, angry music with W2V2 does not show a performance drop, unlike other emotions with W2V2 for either speech or music. We believe that the deep layers of W2V2 might be particularly effective at modeling acoustic characteristics related to angry music, such as intense rhythm. These observations suggest that different emotions in speech and music exhibit distinct acoustic characteristics, highlighting the emotion bias problem in both speech and music SSL models, and strengthening the findings of [10]. Finally, other emotions all have their respective variations on speech or music with speech models or the music model, but the overall patterns remains similar. This further verifies previous findings that shared acoustic representations exist between speech and music [6, 13].
@@ V-B1 Experimental Settings
For the baseline and WS models, we use the AdamW optimizer with a learning rate of $1e{-3}$. For the PEFT models, we use the AdamW optimizer with a learning rate of $1e{-4}$. For the same reason as for the layerwise probing, each model is trained for 300 epochs at each stage to achieve its best possible performance.
@@ V-B2 Results and Discussions
Table III summarizes the domain adaptation results using all three approaches and models. Note that the results of SER and MER from the same model are not comparable, as they are different tasks. We only compare SER with SER and MER with MER. It can be seen that for the speech models (W2V2 and HuBERT), fine-tuning on either domain in stage one always helps the other domain in stage two. For example, the performances of SER followed by MER (74.88, 80.30, 87.20 for W2V2; 84.24, 82.76, 93.10 for HuBERT) are always better than directly conducting SER (73.89, 79.31, 87.20 for W2V2; 83.25, 80.30, 88.18 for HuBERT). Furthermore, PEFT works extremely effectively for HuBERT, achieving the best performance (93.10) and the largest improvement (4.92). For the music model, while several performances are improved after domain adaptation (e.g., 72.77 vs. 61.85 for SER and 93.60 vs. 91.63), such a phenomenon does not generally hold. This finding further verifies our conclusion of Finding III in Section V-A that speech SSL models are further trainable with music acoustics, yet further training of music SSL models with speech acoustics may be less effective.
@@ V-C2 Results and Discussions
(I) In terms of distance ranking, all the speech and music models exhibit consistent patterns: angry and fearful have the smallest distances, followed by happy, sad, neutral, and calm. These consistent patterns indicate that the SSL models have a similar ability to capture acoustic similarities related to emotion in their respective representations. Since speech and music share the same text in RAVDESS, the low FAD scores for certain emotions indicate that their acoustics are largely similar between speech and music. For example, angry speech and music are more acoustically similar than calm speech and music.
@@ VI Conclusion
As the first work exploring acoustic similarity between emotional speech and music through SSL models, we conduct layerwise probing to explore the models’ ability to capture emotion-relevant acoustic cues and uncover the commonality and differences between speech and music SSL models. Moreover, aiming to leverage each domain to enhance the emotion recognition performance for the other domain, we implement and compare three approaches, demonstrating efficacy of the PEFT approach for cross-domain adaptation and noticing the domain-specificity of the models. Finally, we employ FAD to measure the acoustic similarity between speech and music representations on each emotion, revealing different behaviors of SSL models and the emotion-bias issue in both domains. Our study provides new insights into the acoustic similarity between emotional speech and music, particularly with SSL models. We also demonstrate the potential to leverage knowledge from each domain to facilitate the other. As there are limited datasets suitable for further investigation and fair comparison, we acknowledge the limitations of our work and plan to collect new datasets for extending our study in the future.